OpenStudy (rissaunnie):

fan and medal for the solution set t^2-121=0

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

factor as the difference of two squares \[(t+11)(t-11)=0\] then set each factor equal to zero and solve

OpenStudy (misty1212):

or you can add 121 to both sides and get \[t^2=121\\t=\pm\sqrt{121}\\ t=\pm11\]

OpenStudy (jango_in_dtown):

the solution set is {-11,11}

OpenStudy (anonymous):

doing it for the message, the solution set is {-11,11}